if tanA = -(tanB)/2 = -1/3 then find A in terms of  pi.

Grade:11

1 Answers

Vikas TU
7 years ago
 
if tanA = -(tanB)/2 = -1/3
Given tanA = -1/3
A = arctan(-1/3) => 
-0.321 radian
And tanB = 2/3
B = arctan(2/3)
= >  0.588 radian